Suppose T(n) and f(n) and two functions. Write asymptotic notations (Ο, Ω, Θ) using these two functions and explain the growth rate of these functions in each notation.


1
Expert's answer
2021-11-23T13:12:10-0500

"f(n)=O(T(n))" if there exists a positive real number M and a real number n0 such that

"f(n)\le MT(n)" for all "n\ge n_0"


"f(n)=\Omega(T(n))" if there exists a positive real number M and a real number n0 such that

"f(n)\ge MT(n)" for all "n\ge n_0"


"f(n)=\theta(T(n))" if there exists a positive real numbers M1, M2 and a real number n0 such that

"M_1T(n)\le f(n)\le M_2T(n)" for all "n\ge n_0"